OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. c times the square root of c + 12 times c times the square root of c − 2 times c times the square root of c 11c c times the square root of c 11 times c times the square root of c 11 the square root of c

OpenStudy (phi):

can you use the equation editor ?

OpenStudy (anonymous):

How would I do that? lol

OpenStudy (anonymous):

nvrm lol

OpenStudy (phi):

use the button in the lower left

OpenStudy (anonymous):

I don't know how to do this....I can try to draw it?

OpenStudy (phi):

\[c \sqrt{c}+ 12c \sqrt{c}-2c \sqrt{c}\]

OpenStudy (anonymous):



OpenStudy (phi):

try factoring out c sqrt(c) from each term what do you get ?

OpenStudy (anonymous):

one sec, I'm a visual learner so I write it all down

OpenStudy (anonymous):

would you combine all of the c's?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

factor means "undo distribution" if you had (x + 2x + 3x) and factor out x you get x(1+2+3) if you "distribute" the x, you get x+2x+3x which is what you started with another example: (x y + 3 xy) factor out x and y: xy(1 + 3)

OpenStudy (anonymous):